One thousand kilograms per hour of a mixture of benzene (B) and toluene (T) that contains 60 % benzene by mass are separated by distillation into two fractions. The mass flow rate of benzene in the top stream is 470 kg B/h, and that of toluene in the bottom stream is 350 kg T/h. The operation is at steady state. a. Draw and label the flowchart of the process. b. Write balances on benzene and toluene to calculate the unknown component flow rates in the output streams.


1
Expert's answer
2021-11-13T02:09:48-0500



Overall balance 

"1000=D+B.........(1)"

Applying benzene balance 

"(1000)(0.5)=D \times D + B \times B = 450 + B \times B\\
B \times B =50"

This is the benzene flow rate at the bottom


Applying tolere balance 

"(1000)(0.5)=D y D + B y B = DyD+475\\
DyD= 25"

This is the tolere flow rate at the top